# A Fast and Optimal Learning-based Path Planning Method for Planetary Rovers

### v2

@@ Abstract
Intelligent autonomous path planning is crucial to improve the exploration efficiency of planetary rovers. In this paper, we propose a learning-based method to quickly search for optimal paths in an elevation map, which is called NNPP. The NNPP model learns semantic information about start and goal locations, as well as map representations, from numerous pre-annotated optimal path demonstrations, and produces a probabilistic distribution over each pixel representing the likelihood of it belonging to an optimal path on the map. More specifically, the paper computes the traversal cost for each grid cell from the slope, roughness and elevation difference obtained from the DEM. Subsequently, the start and goal locations are encoded using a Gaussian distribution and different location encoding parameters are analyzed for their effect on model performance. After training, the NNPP model is able to perform path planning on novel maps. Experiments show that the guidance field generated by the NNPP model can significantly reduce the search time for optimal paths under the same hardware conditions, and the advantage of NNPP increases with the scale of the map.
@@
path planning, deep learning, planetary rovers, traversability cost, elevation map
@@
[1,2,3,4,5,6]organization=State Key Laboratory of Robotics and Systems, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, state=Heilongjiang, country=China
@@ 1 INTRODUCTION
During planetary exploration missions, traversing rough terrain replete with obstacles poses challenges to rovers tasked with exploring diverse targets of interests. The surfaces of extraterrestrial bodies such as Mars and the Moon are beset with hazardous features including rocks, steep slopes and craters, while rovers are constrained by limited onboard energy reservessutoh2015right . Hence, identifying collision-free and energy-efficient path to destination sites emerges as a crucial requirement for planetary rovers aimed at extending their operational lifetimes and scientific output. Moreover, planetary rovers encounter challenging navigation issues during both short and long-range planning stages. The navigation problems are further exacerbated by the limited bandwidth and time lag for communication with Earth, which restricts the rovers to only one round-trip communication per sol (a Martian day that lasts approximately 24 hours and 40 minutes) in operational terms. Currently, the existing technology for driving Mars rovers mainly relies on images captured from the ground to plan movements. However, these plans can only be executed within the range visible to the rover. This underscores the importance of designing autonomous path planning algorithms for planetary rovers.
The planetary path planning problem can be classified into global path planning and local path planning. Global path planning requires determining the entire path from the rover’s starting position to its intended target using planetary surface images captured by orbiting satellites. On the other hand, local path planning is based on online maps generated by onboard sensors, making it suitable for unknown and dynamic environments. This paper focuses on addressing the global path planning problem for planetary rovers.
@@
To address the navigation problem of mobile robots in rugged terrains such as planet surfaces, researchers initially proposed the idea of binary terrain classification into impassable obstacles and free traversal regionslaubach1998autonomous . In ettlin2006rough , Ettlin suggested combining binary motion planners with traversability classification by integrating rapidly-exploring random trees (RRT) and gradually increasing the threshold for binary terrain classification. The previous methods may be too conservative since classifying an obstacle as non-traversable can result in a suboptimal plan or even failure to find a path. Given the rough nature of the terrain, an obstacle may be traversable from a particular direction or at a specific velocity. The issue of environment representation is of paramount importance in the domain of rough terrain navigation for autonomous systems. Hence, an adequate model that captures the accessibility of the terrain is needed. In traditional binary approaches, the environment is binarized into traversable and non-traversable categories. However, in rough terrain navigation it is not clear how to binarize the environment in a meaningful and robust manner, given the continuum of terrain difficulties. Therefore, a continuous approach that assigns a continuous cost value for traversing different parts of the terrain should be considered. Such an approach would provide a more nuanced representation of the environment, thereby allowing for more informed navigation planning.
The accuracy of global path planning depends on the scene information contained within the map. This information includes the layout of the terrain, the presence of obstacles, environmental conditions, and other relevant factors. These elements are crucial for a successful plan as they contribute to the understanding of the environment, which is essential for determining an optimal route. Compared to RGB images obtained by orbiting satellites, using a stereo camera CCD to capture stereo image pairs in orbit and generating digital elevation model (DEM) data containing terrain height information can provide more accurate elevation data of the planetary surface. In uneven planetary terrain, map data that includes height information can better reflect the requirements of planetary rovers for terrain exploration.
@@
Thus, this paper adopts DEM data as the map representation and formulates the global path planning problem as finding the optimal feasible path that avoids untraversable areas, including obstacles, and minimizes the risk of failure due to hazardous terrain morphology, as shown in Figure. 1.
@@
The remainder of the paper is organized as follows. Section 2 introduces related work on path planning, focusing on learning-based path planning methods. Section 3 presents environment modeling approaches and cost map computation methods, where the path with the lowest total cost is identified as the optimal path. Section 4 elaborates on the proposed method, including the network structure, dataset preparation, and model evaluation metrics. Section 5 introduces the simulation experiments designed for the model, which demonstrates that the model can significantly reduce optimal path search time.
@@ 2 RELATED WORKS
Numerous algorithms have been proposed and refined to tackle the path planning problem, including the ${A^{\star}}$ algorithmhart1968formal , RRT-based algorithmslavalle2001randomized , artificial potential field (APF) based algorithmsniu2023virtual , and intelligent algorithms such as the genetic algorithmzhang2016dynamic , ant colony algorithmliu2017improved and reinforcement learning based algorithmsjaradat2011reinforcement . This section focuses on the mainstream methods of path planning and their applications in rough terrain navigation.
@@ 2.1 Heuristic method
Heuristic path planning algorithms utilize a heuristic function ${H\left(n\right)}$ to select the extended nodes, significantly reducing the total number of extended nodes and improving the efficiency of trajectory planningwang2021warehouse . In a Grid map environment, the ${A^{\star}}$ algorithm can effectively represent obstacles of arbitrary shapes in two-dimensional space. However, in a three-dimensional environment, the evaluation function becomes complicated with the increase in dimensionality. Thus, the ${A^{\star}}$ algorithm is less commonly used in high-dimensional space. To address this limitation, many studies have proposed optimization schemes based on ${A^{\star}}$ algorithmsastar2020 , such as ${D^{\star}}$, ${FieldD^{\star}}$, ${Theta^{\star}}$, Anytime Repairing ${A^{\star}}$ (${ARA^{\star}}$), and more.
In the elevation map scenario, Bo Zhangzhang2022path utilized the cumulative traversability cost between grids as the cost function for ${A^{\star}}$ algorithm, thereby enabling it to search for the shortest path in 3D uneven terrains.
@@ 2.2 Learning based methods
In recent years, modulated by the rapid development and widespread applications of artificial intelligence technologies, especially the advances within deep reinforcement learning (DRL), learning-based methods for path planning, obstacle detection, trafficability analysis and related path reasoning tasks have received increasing interest from researchers in both academia and industry hoping to enable more autonomous navigation for unmanned systems operating in complex environmentstai2017virtual zhou2019vision .
@@
Based on the heuristic methods discussed in the previous section, many researchers have employed learning-based techniques to construct guiding functions that accelerate the convergence process of classical algorithms. This approach has proven to be highly effective. The ${NRRT^{\star}}$wang2020neural approach applies a convolutional neural network (CNN) model to guide the sampling-based RRT path search on a 2D grid map while satisfying different clearance and step size planning requirements. The same team of authors proposed the NEED modelwang2021robot , which predicts promising search regions to guide the search direction of planning algorithms. This approach significantly enhances algorithm performance on new problems while maintaining a prediction speed of over 100 Hz on a laptop for high-dynamics applications. Reinforcement learning (RL), compared to traditional control methods, has the notable advantages of not requiring environmental maps, possessing strong learning capabilities from data, and exhibiting high dynamic adaptability to changes in complex environmentsmoreira2020deep . Its deep neural network architecture can effectively process high-dimensional information from various sensors, while its reinforcement learning mechanism can perform continuous decision-making tasks in complex and dynamic environments. This makes it particularly suitable for local path planning in dynamic scenarios. Many researchers have attempted to apply reinforcement learning methods to local path planning in rough terrain. Q-learning has been applied to local path planning, and according to Reference zhang2022path , reinforcement learning methods can yield superior paths compared to traditional approaches. Reference josef2020deep indicates that compared to baseline methods, reinforcement learning techniques can improve the success rate of path planning. It can also be inferred that reinforcement learning methods consume more time compared to traditional approaches.
@@
The time taken by the Q-learning method is almost double that of the APF method and approximately 15 times that of the conventional ${A^{\star}}$ method. The success rate of reinforcement learning (95%-96%) also impacts its applicability to global maps at a larger scale. For example, Reference zhang2022path reports that the proposed reinforcement learning method consumes nearly 11 times the time compared to traditional ${A^{\star}}$ methods when searching paths in a ${110\times 110}$ elevation map, while only improving path optimality by 20%. The proposed NNPP method in this paper not only greatly reduces the search time (shortened by 4.6 times) and algorithm optimality is nearly on par with ${A^{\star}}$ (reduced by 2%) compared to traditional ${A^{\star}}$ on the same hardware and a ${256\times 256}$ map, but NNPP’s performance advantage will further increase with the expansion of the map size.
@@ 3 ENVIRONMENTAL MODELING AND PROBLEM FORMULATION
Different environmental modeling approaches can have different effects on the optimal path search results. DEM is commonly used in many instances of path planning on uneven terrains, thus this chapter introduces DEM-based environmental modeling approaches as well as the problem formulation discussed in this paper.
@@ 3.1 Digital Representation of Elevation Map
In practical applications of CE4wang2020vision , digital elevation model (DEM) is used for map generation and path searching. The spatial distribution of an uneven terrain environment is characterized utilizing a group of three-dimensional vectors ${\left\{{x,y,z}\right\}}$ in which ${x}$ and ${y}$ represent the planar coordinates of mapping an environmental feature point within the uneven terrain while ${z}$ denotes the altitude of the point. Therefore, geomorphic attributes including aspect, gradient of slope, changes in elevation, and roughness can be adequately extracted via DEM.
@@ 3.2 Information Extraction of Uneven Terrain Environment
According to the traversability assessment method used by Yutu 2 lunar roveryan2019autonomous , geomorphic attributes of the ${i{\rm{-th}}}$ grid cell in DEM data are calculated by the Eq. 1. To calculate the roughness, a patch consisting of the ${i{\rm{-th}}}$ grid cell and its eight neighboring cells is used.
@@
The angle between the normal vector and horizontal plane of grid ${i}$ is defined as ${\varphi_{i}}$. The variable ${r}$ denotes the residual of the fitting distance from each DEM data point in a unit cell to the fitted plane in the corresponding patch, representing the roughness of the terrain region. The variable ${\Delta h}$ represents the height difference between each DEM data point in a unit cell and the fitted plane in the corresponding patch, representing the elevation difference in the terrain region. The traversability cost for each grid cell is calculated as shown in Figure. 9. In Eq. 1, ${\varphi_{s}}$ represents the maximum safe tilt angle that a rover can traverse through a single grid cell, ${H_{s}}$ represents the maximum obstacle height value, and rs represents the maximum roughness value that a rover can navigate through. The variables ${k_{s}}$, ${k_{r}}$, and ${k_{f}}$ are weighting coefficients, and their sum is equal to 1. Let the traversability cost for traveling between the two grids ${i}$ and ${j}$ be Eq. 2.
${T_{i,j}}=\sqrt{{{\left({{T_{i}}-{T_{j}}}\right)}^{2}}}$ (2)
@@
We adopt the approach of modeling the environment by calculating the cost of passing through each pixel value because in rugged and uneven terrains, the shortest Euclidean path may be impassable, thus it is necessary to integrally describe the terrain environment by combining slope, roughness and other geometric features in order to find the optimal traversable path. After the aforementioned environmental modeling, the problem discussed in this paper can be abstracted as searching for a path from a given start point to end point in a 2.5D cost map such that the path has the minimum accumulated traversability cost.

### v3

# NNPP: A Learning-Based Heuristic Model for Accelerating Optimal Path Planning on Uneven Terrain

@@ Abstract
Intelligent autonomous path planning is essential for enhancing the exploration efficiency of mobile robots operating in uneven terrains like planetary surfaces and off-road environments. In this paper, we propose the NNPP model for computing the heuristic region, enabling foundation algorithms like ${A^{\star}}$ to find the optimal path solely within this reduced search space, effectively decreasing the search time. The NNPP model learns semantic information about start and goal locations, as well as map representations, from numerous pre-annotated optimal path demonstrations, and produces a probabilistic distribution over each pixel representing the likelihood of it belonging to an optimal path on the map. More specifically, the paper computes the traversal cost for each grid cell from the slope, roughness and elevation difference obtained from the digital elevation model. Subsequently, the start and goal locations are encoded using a Gaussian distribution and different location encoding parameters are analyzed for their effect on model performance. After training, the NNPP model is able to accelerate path planning on novel maps. Experiments demonstrate that the heuristic region generated by the NNPP model achieves a 3$\times$×speedup for optimal path planning under identical hardware conditions. Moreover, the NNPP model’s advantage becomes more pronounced as the size of the map increases.
@@
path planning , deep learning , autonomous navigation , elevation map
journal: Arxiv
@@
[1,2,3,4,5,6]organization=State Key Laboratory of Robotics and System, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, state=Heilongjiang, country=China
@@ 1 INTRODUCTION
Path planning is the process of finding a desired path between two given locations while satisfying predefined requirements. With the emergence of field mobile robotics, such as rescue robots [1] and planetary rovers [2], path planning in outdoor environments has gained significant attention recently. Outdoor unstructured environments often feature uneven terrains where some paths are too hazardous for mobile robots to traverse due to instability on steep slopes and limited motion power [3]. Furthermore, unlike indoor scenes, the shortest paths in uneven terrains are not always optimal for mobile platforms as they often have to deal with rapid elevation changes. When mapping routes through rugged terrains, it’s essential to account for more than just obstacles such as jutting rocks and sudden drops. It is also necessary to strike a balance between the shortest geometric distance and the potential dangers posed by uneven terrain. This complexity often stumps conventional algorithms like ${A^{\star}}$ and RRT, either leading to suboptimal paths or excessive computation time [4]. Therefore, modeling terrain and conducting traversability analysis, as well as establishing criteria for evaluating the quality of a path, are essential prerequisites for path planning.
@@
To address the traversability analysis problem of mobile robots in rugged terrains such as planet surfaces, researchers initially proposed the idea of binary terrain classification into impassable obstacles and free traversal regions [5]. In [6], Ettlin suggested combining binary motion planners with traversability classification by integrating rapidly-exploring random trees and gradually increasing the threshold for binary terrain classification. However, when navigating rough terrain, this binary categorization proves ineffective due to the continuous spectrum of terrain complexities since classifying an obstacle as non-traversable can result in a suboptimal plan or even failure to find a path. Thus, an adequate model capturing terrain traversability is essential for a nuanced environment representation and informed navigation planning.
The traversability of the uneven terrain depends on the scene information contained within the map. This information includes the layout of the terrain, the presence of obstacles, environmental conditions, and other relevant factors. For instances, in uneven planetary terrain, map data that includes height information can better reflect the requirements of planetary rovers for terrain exploration.
Figure 1: In the proposed framework, a cost map with start and goal points is fed into the NNPP. The model generates a probability map where each pixel’s probability value indicates the likelihood of the optimal path passing through that pixel. This probability map is then utilized to guide the ${A^{\star}}$ algorithm on the original map, resulting in the final optimal path.
@@
Thus, this paper adopts DEM data as the uneven map representation and formulates the global path planning problem as finding the optimal feasible path that avoids untraversable areas, including obstacles, and minimizes the risk of failure due to hazardous terrain morphology, as shown in Figure 1.
@@
The remainder of the paper is organized as follows. Section 2 introduces related work on path planning, focusing on learning-based path planning methods. Section 3 presents environmental modeling and cost map calculation methods, and investigates the trade-off between minimizing total cost and minimizing total path length. Section 4 elaborates on the proposed method, including the network structure, dataset preparation, and model evaluation metrics. Section 5 introduces the simulation experiments designed for the model, which demonstrates that the model can significantly reduce optimal path search time.
@@ 2 RELATED WORKS
Numerous algorithms have been proposed and refined to tackle the path planning problem, including the ${A^{\star}}$ algorithm [7], RRT-based algorithms [8], artificial potential field (APF) based algorithms [9], and intelligent algorithms such as the genetic algorithm [10], ant colony algorithm [11] and reinforcement learning based algorithms [12]. This section focuses on the mainstream methods of path planning and their applications in rough terrain navigation.
@@ 2.1 Heuristic method
Heuristic path planning algorithms utilize a heuristic function ${H\left(n\right)}$ to select the extended nodes, significantly reducing the total number of extended nodes and improving the efficiency of trajectory planning [13]. In a Grid map environment, the ${A^{\star}}$ algorithm can effectively represent obstacles of arbitrary shapes in two-dimensional space. However, in a three-dimensional environment, the evaluation function becomes complicated with the increase in dimensionality. Thus, the ${A^{\star}}$ algorithm is less commonly used in high-dimensional space. To address this limitation, many studies have proposed optimization schemes based on ${A^{\star}}$ algorithms [14], such as ${D^{\star}}$, ${FieldD^{\star}}$, ${Theta^{\star}}$, Anytime Repairing ${A^{\star}}$ (${ARA^{\star}}$), and more.
In the elevation map scenario, Bo Zhang [15] utilized the cumulative traversability cost between grids as the cost function for ${A^{\star}}$ algorithm, thereby enabling it to search for the shortest path in 3D uneven terrains.
@@ 2.2 Learning based methods
In recent years, modulated by the rapid development and widespread applications of artificial intelligence technologies, especially the advances within deep reinforcement learning (DRL), learning-based methods for path planning, obstacle detection, trafficability analysis and related path reasoning tasks have received increasing interest from researchers in both academia and industry hoping to enable more autonomous navigation for unmanned systems operating in complex environments [16] [17].
@@
Based on the heuristic methods discussed in the previous section, many researchers have employed learning-based techniques to construct guiding functions that accelerate the convergence process of classical algorithms. This approach has proven to be highly effective. The ${NRRT^{\star}}$ [18] approach applies a convolutional neural network (CNN) model to guide the sampling-based RRT path search on a 2D grid map while satisfying different clearance and step size planning requirements. The same team of authors proposed the NEED model [19], which predicts promising search regions to guide the search direction of planning algorithms. This approach significantly enhances algorithm performance on new problems while maintaining a prediction speed of over 100 Hz on a laptop for high-dynamics applications. Reinforcement learning (RL), compared to traditional control methods, has the notable advantages of not requiring environmental maps, possessing strong learning capabilities from data, and exhibiting high dynamic adaptability to changes in complex environments [20]. Its deep neural network architecture can effectively process high-dimensional information from various sensors, while its reinforcement learning mechanism can perform continuous decision-making tasks in complex and dynamic environments. This makes it particularly suitable for local path planning in dynamic scenarios. Many researchers have attempted to apply reinforcement learning methods to local path planning in rough terrain. Q-learning has been applied to local path planning, and according to Reference [15], reinforcement learning methods can yield superior paths compared to traditional approaches. Reference [21] indicates that compared to baseline methods, reinforcement learning techniques can improve the success rate of path planning. It can also be inferred that reinforcement learning methods consume more time compared to traditional approaches.
@@
The time taken by the Q-learning method is almost double that of the APF method and approximately 15 times that of the conventional ${A^{\star}}$ method. The success rate of reinforcement learning (95%-96%) also impacts its applicability to global maps at a larger scale. For example, Reference [15] reports that the proposed reinforcement learning method consumes nearly 11 times the time compared to traditional ${A^{\star}}$ methods when searching paths in a ${110\times 110}$ elevation map, while only improving path optimality by 20%. The proposed NNPP method in this paper not only greatly reduces the search time (shortened by 3 times) and algorithm optimality is nearly on par with ${A^{\star}}$ (reduced by 2%) compared to traditional ${A^{\star}}$ on the same hardware and a ${256\times 256}$ map, but NNPP’s performance advantage will further increase with the expansion of the map size.
@@ 3 ENVIRONMENTAL MODELING AND PROBLEM FORMULATION
The environmental modeling processes, such as grid-based, graph-based, and point-based methods, are prerequisites for path planning. Digital Elevation Model (DEM) is commonly used in many instances of path planning on uneven terrains, thus this chapter introduces DEM-based environmental modeling approaches as well as the problem formulation discussed in this paper.
@@ 3.1 Digital Representation of Elevation Map
In practical applications of CE4 [22], DEM is used for map generation and path searching. The spatial distribution of an uneven terrain environment is characterized utilizing a group of three-dimensional vectors ${\left\{{x,y,z}\right\}}$ in which ${x}$ and ${y}$ represent the planar coordinates of mapping an environmental feature point within the uneven terrain while ${z}$ denotes the altitude of the point. Therefore, geomorphic attributes including aspect, gradient of slope, changes in elevation, and roughness can be adequately extracted via DEM.
@@ 3.2 Information Extraction of Uneven Terrain Environment
According to the traversability assessment method used by Yutu 2 lunar rover [23], geomorphic attributes of the ${i{\rm{-th}}}$ grid cell in DEM data are calculated by the Eq.1. To calculate the roughness, a patch consisting of the ${i{\rm{-th}}}$ grid cell and its eight neighboring cells is used.
@@
The angle between the normal vector and horizontal plane of grid ${i}$ is defined as ${\varphi_{i}}$. The variable ${r}$ denotes the residual of the fitting distance from each DEM data point in a unit cell to the fitted plane in the corresponding patch, representing the roughness of the terrain region. The variable ${\Delta h}$ represents the height difference between each DEM data point in a unit cell and the fitted plane in the corresponding patch, representing the elevation difference in the terrain region. The traversability cost for each grid cell is calculated as shown in Figure 2.
Figure 2: The traversal cost for each pixel was calculated using the following process: ${(a)}$ the original DEM was obtained; ${(b)}$ the slope cost was computed; ${(c)}$ the roughness cost was calculated; ${(d)}$ the elevation difference cost was determined. The resulting images, displayed from left to right, represent the original DEM, slope cost, roughness cost, and elevation difference cost. In this paper, we set ${\varphi_{s}=30\textdegree}$, ${H_{s}=0.2}$, ${r_{s}=0.6}$.
In Eq.1, ${\varphi_{s}}$ represents the maximum safe tilt angle that a rover can traverse through a single grid cell, ${H_{s}}$ represents the maximum obstacle height value, and ${r_{s}}$ represents the maximum roughness value that a rover can navigate through. The variables ${k_{s}}$, ${k_{r}}$, and ${k_{f}}$ are weighting coefficients, and their sum is equal to 1. In this study, we set the parameters ${k_{s}}$, ${k_{r}}$, and ${k_{f}}$ to 0.6, 0.2, and 0.2, respectively. Among these parameters, ${k_{s}}$ has the greatest impact on the robot’s performance, as it accounts for the influence of terrain slope. Therefore, we assign a larger value to ${k_{s}}$ compared to ${k_{r}}$ and ${k_{f}}$ [24]. The specific values of ${\varphi_{s}}$, ${H_{s}}$ and ${r_{s}}$ are determined by the robot’s structural and kinematic-dynamic parameters. As this paper focuses on planetary surfaces as a research case, we have referenced relevant data from lunar and Martian rovers [25] to establish values for ${\varphi_{s}}$, ${H_{s}}$ and ${r_{s}}$ set respectively at 30 deg, 0.2 m, and 0.6. Let the traversability cost for traveling between the two grids $i$ and ${i+1}$ be Eq.2.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{T_{i,i+1}}=T_{i}+T_{i+1}}$ (2)
@@
We adopt the approach of modeling the environment by calculating the cost of passing through each pixel value because in rugged and uneven terrains, the shortest Euclidean path may be impassable, thus it is necessary to integrally describe the terrain environment by combining slope, roughness and other geometric features in order to find the optimal traversable path. After the aforementioned environmental modeling, the problem discussed in this paper can be abstracted as searching for a path from a given start point to the end in a 2.5D cost map that strikes an ideal balance between accumulated traversability cost and Euclidean path length.
3.3 Task definition
As described in previous sections, the DEM of the uneven terrain is transformed into a traversability cost map as shown in Figure 3. Each grid cell in the map is assigned a cost value between 0 and 1. A path is composed of a sequence of adjacent grid coordinates. The Euclidean length of a path is the sum of the Euclidean distances between adjacent grid cells along the path, as shown in Eq.3. The cumulative traverse cost of a path represents the sum of the traverse cost values of the grid cells ($T_{i}$) along the path, denoted as $CC$, as Eq.4.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{L=\sum{{}_{i=1}^{n-1}}{\sqrt{(X_{i+1}-X_{i})^{2}+(Y_{i+1}-Y_{i})^{2}}}}}$ (3)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{{CC}=\sum{{}_{i=0}^{n}}{{T_{i}}}}}$ (4)
In path planning tasks, the input consists of a traversability cost map and the start and goal coordinates, and the model outputs a path represented by a sequence of grid coordinates. This path minimizes the weighted length (${L}+\omega\cdot{CC}$). The intricate relationship between the $\omega$ and the balance between $CC$ and $L$ is delved into in Section 4.2, leading to the identification of an optimal value of $\omega$.
Figure 3: Path planning task, including a traversability cost map illustration and a path illustration